OpenStudy (anonymous):

Solve the proportion 2 times x all over 6 times x plus 7 equals 2 over x. What is the value of x?

OpenStudy (anonymous):

\[2x/(6x +7) =2/x\]

OpenStudy (anonymous):

Like this?

OpenStudy (anonymous):

Yes. So would you cross multiply 2(6x+7) then divide by 2x? I am kinda lost. :( the possible answers are a. x=1 and x=-7 b. x=-1 and x=7 c. x=-1 and x=6 d. x=1 and x=6

OpenStudy (anonymous):

Actually since they equal each other then if you subtract one from the other you would get 0

OpenStudy (anonymous):

\[(2x/(6x+7 ) - 2/x = 0\]

OpenStudy (anonymous):

Then you need to find the least common denominator and subtract..what do you come up with?

OpenStudy (anonymous):

2x?

OpenStudy (anonymous):

(6x+7)(x) is the least common denominator. Multiply it back into your two numbers and what equation do you get?

OpenStudy (anonymous):

So the least common denominator is x or (6x+7)(x)?

OpenStudy (anonymous):

I am sorry I am having a really hard time with this problem.

OpenStudy (anonymous):

(6x+7)(x)..do you see how a got that. When dealing with fraction that ust be added or subtracted they must have the same denominator. Most all the time the LCD (Least common denominator is multiplying the two denominators together. Now you must multiply each number by the LCD. What will your equation look like now?

OpenStudy (anonymous):

\[ \frac{ 2x }{ 6x+7} - \frac{ 2 }{ x } = 0\] Do you see how we got this?

OpenStudy (anonymous):

Yes I got it so far

OpenStudy (anonymous):

You subtracted 2/x from both sides to get a 0 on the right and a -2/x on the left

OpenStudy (anonymous):



OpenStudy (anonymous):

If a =b then a-b will always be 0.

OpenStudy (anonymous):

Can you solve these

OpenStudy (anonymous):

Ok take your time

OpenStudy (anonymous):

Do any of these cancel out?

OpenStudy (anonymous):

Yes very good..go on

OpenStudy (anonymous):

so the top one would be 2x/x because the 6x+7 cancel right? And the bottom would be 12x+14 and the x would cancel? Am I even close?

OpenStudy (anonymous):

The top one is (2x)(x) which equals what? The bottom one is correct very good

OpenStudy (anonymous):

\[2x ^{2}\]

OpenStudy (anonymous):

Now what do you do?

OpenStudy (anonymous):

reduce?

OpenStudy (anonymous):

Remember we were subtracting one from the other to get 0. So we subtract 2x^2 -(12x+14) and this all needs to be over our LCD. So what do we have?

OpenStudy (anonymous):

NO idea, I am lost now.

OpenStudy (anonymous):

What would this look like when you subtract?\[2x ^{2} -(12x+14)\]

OpenStudy (anonymous):

Factor it? \[2(x ^{2}-6x+7)\]

OpenStudy (anonymous):

Close...you a\[2x ^{2}-12x-14\]re a step ahead of me...The equation would look like this

OpenStudy (anonymous):

You see how I got that?

OpenStudy (anonymous):

ok so then it would be a -7 instead of a +7 like I put it

OpenStudy (anonymous):

That's right

OpenStudy (anonymous):

2(x-7)(x+1)

OpenStudy (anonymous):

So our equation will look like this. \[\frac{ 2x ^{2}-12x-14 }{ (6x+7)(x)} = 0\]

OpenStudy (anonymous):

ok Not sure what to do next

OpenStudy (anonymous):

Factored it would look like this \[\frac{ (2)(x-7)(x+1) }{ (6x+7)(x) } =0\]

OpenStudy (anonymous):

Do you see this?

OpenStudy (anonymous):

Okay that is what I did above. Yay me!

OpenStudy (anonymous):

You are doing great...Now for a fraction to equal 0 what would it look like?

OpenStudy (anonymous):

Still not sure where to go from here The bottom part of the fraction is throwing me off

OpenStudy (anonymous):

The bottom is unimportant since we are only looking to find solutions that equal zero..

OpenStudy (anonymous):

No idea what a fraction equal to 0 would look like

OpenStudy (anonymous):

0 = 0

OpenStudy (anonymous):

0/a =0, where a = any number This means zero over any number will always equal 0 0/4 = 0 0/12 - 0 0/5000 =0

OpenStudy (anonymous):

Still lost

OpenStudy (anonymous):

What are we trying to get our factored fraction to equal??

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

Right..so to do this we have to get the numerator (The top of the fraction to equal 0) because 0 over any number equals 0. Do you see where I am going?

OpenStudy (anonymous):

No I don't see where you are going. I am sorry. :(

OpenStudy (anonymous):

Not a problem..is it the concept of 0/4 or 0/8 you do not understand?

OpenStudy (anonymous):

I understand that both those = 0. I just don't understand the next steps and what we are trying to do to find what x equals.

OpenStudy (anonymous):

Ok so we now have to get the numerator to equal 0. Our factored numerator is \[(2)(x-7)(x+1)\]

OpenStudy (anonymous):

How or what can we make the values of x to make this equal 0?

OpenStudy (anonymous):

The opposite of what they already are?

OpenStudy (anonymous):

What happens if I make x=7?

OpenStudy (anonymous):

How did you make x=7?

OpenStudy (anonymous):

Remember we are trying to make the numerator = 0. Our numerator is (2)(x-7)(x+1) So we are trying to solve the following equation.. 2(x-7)(x+1) = 0. How would you solve for x?

OpenStudy (anonymous):

Remember 0 multiplied by any number always equals zero. 5 x 0 = 0 6 x 0 = 0 7 x 5 x 0 = 0

OpenStudy (anonymous):

No idea

OpenStudy (anonymous):

If (2) times (x-7) times (x+1) = 0....What happens when I make x =7? It becomes (2) times (7-7) times (7+1) =0 Or (2) times 0 times 8 which equals??

OpenStudy (anonymous):

0

OpenStudy (anonymous):

So the answer must be b because that is the only one with a 7 in it.

OpenStudy (anonymous):

Right,,When dealing with quadratic equations that equal 0. All you need to do is solve for each value of x so that the factor is 0. Your two factors are (x-7) and (x+1). So once you solve for these two unkowns you have your answer.

OpenStudy (anonymous):

Your answers is 7 and -1.

OpenStudy (anonymous):

Trust me this get easier with practice...

OpenStudy (anonymous):

Thank you, I am sure I will be back soon. :)

OpenStudy (anonymous):

You did well..

OpenStudy (anonymous):

LOL! you are too nice. An hour later and I barely understand it. Thanks for being so patient.